The body was released to fall freely from a height of 100 m above the Earth's surface. Determine its speed just before hitting the ground. When considering this problem ignore air resistance.


1
Expert's answer
2021-03-21T11:24:05-0400
"PE=KE,""mgh=\dfrac{1}{2}mv^2,""v=\sqrt{2gh}=\sqrt{2\cdot9.8\ \dfrac{m}{s^2}\cdot100\ m}=44.27\ \dfrac{m}{s}."